# Bayesian Multifractal Image Segmentation

### v1

@@ 6.4.2 Detection Performance
Figure 4 shows the analyzed section of (a) the mammography image with (b) its respective segmentation mask, highlighting the tumor, and (c-j) the segmentation results of the proposed method compared to the baseline methods. The proposed method yields very satisfactory results, with the second best error (0.68%) after the SAM method (0.42%). Note that SAM produces more false negatives in the region of interest (FN: $203$) than the proposed method (FN: $53$). Also note that SAM operates in a supervised manner and requires knowledge of input pixels from each region. In contrast, the proposed method operates in a fully unsupervised mode, demonstrating its potential effectiveness in modeling and handling tumor segmentation tasks.
@@ 7 Conclusions
This work introduced a new Bayesian multifractal image segmentation method formulated as the joint estimation of the multifractality parameters and pixel labels. As a novel key ingredient, the proposed approach leverages on a flexible Fourier domain model for log-leaders allowing the multifractality estimation of heterogeneous images. To model the space-scale label correlations, a multiscale graph structure was introduced and used in a Potts Markov random field prior. A Bayesian formulation was finally introduced to estimate the unknown model parameters (including the multifractality of each region of the image, the image labels and the granularity parameters of the image regions), numerically and efficiently computed using samples generated by an appropriate Gibbs sampler. Numerical experiments conducted on synthetic multifractal images and real data demonstrate the excellent performance of the proposed approach for image segmentation, outperforming several baselines including local multifractality-based approaches, classical texture segmentation methods as well as recent deep learning strategies. Future work includes the extension of the model to data with more than two dimensions (such as depth or spectral data volumes [44]), in which multidimensional/multivariate correlations can be included and exploited in the model. Another potential direction is to evaluate the proposed approach in scenarios with corrupted spatial information, such as noise, missing, or compressed data (as in [45]).

### v2

@@ 6.4.1 Tumor segmentation
The mammography image used for this experiment was extracted from the Mini-DDSM (Digital Database of Screening Mammography) dataset [^5] Mammography dataset available at https://ardisdataset.github.io/MiniDDSM/. The set used for the simulation is “cancer/1727”.. In this dataset, each image has a segmentation label, or contour, that delineates the margins of the tumor [^6] It is important to note that the provided contour serves just as a reference to localize the tumor, as literature has reported that such contours may exhibit inaccuracies. Nevertheless, they offer a sufficiently accurate approximation for characterizing the tumor. The segmentation mask is computed from the provided contour by performing a flood-fill operation on the binary mask.. In this context, the proposed method can be considered with $K=2$ assuming that tumor and normal tissue are characterized by different multifractal parameters. The mammography image was cropped to retain only the region of interest shown in Figure 6-(a), removing the surrounding background. Figure 6 shows the analyzed section of (a) the mammography image with (b) its respective segmentation mask, highlighting the tumor, and (c-j) the segmentation results of the proposed method compared to the baseline methods. The proposed method yields very satisfactory results, with the second best error (0.68%) after the SAM method (0.42%). Note that SAM produces more false negatives in the region of interest (FN: $203$) than the proposed method (FN: $53$). Also note that SAM operates in a supervised manner and requires knowledge of input pixels from each region. In contrast, the proposed method operates in a fully unsupervised mode, demonstrating its potential effectiveness in modeling and handling tumor segmentation tasks.
6.4.2 Water detection from SAR images
Two SAR images were extracted from the dataset in [kim2021large], where the true label for the water segmentation is provided (we considered $K=2$). Figure 7 presents a comparison of segmentation performance, showing that the proposed unsupervised method (c) achieves competitive results and is second best after the FSeg (h) and SAM methods (j): SAM, which is supervised (two points per class were provided as input), yields better performance for the second image but significantly worse for the first, where it fails to detect one of the two water bodies. The FSeg method shows outstanding performance for the first image, but fails quite drastically for the second one. In contrast, the proposed method well detects the water bodies in all cases and yields very satisfactory and robust performance. These results underline the robustness and versatility of the proposed method to handle different segmentation scenarios on challenging images such as SAR data.
Figure 7: Results on SAR images for segmenting water bodies. First and second best performances are highlighted in bold and with an underline, respectively. For SAM, two points per class were provided as prompts to guide the segmentation. All the other methods are unsupervised.
@@ 7 Conclusions
This work introduced a new Bayesian multifractal image segmentation method formulated as the joint estimation of the multifractality parameters and pixel labels. As a novel key ingredient, the proposed approach leverages on a flexible Fourier domain model for log-leaders allowing the multifractality estimation of heterogeneous images. To model the space-scale label correlations, a multiscale graph structure was introduced and used in a Potts Markov random field prior. A Bayesian formulation was finally introduced to estimate the unknown model parameters (including the multifractality of each region of the image, the image labels and the granularity parameters of the image regions), numerically and efficiently computed using samples generated by an appropriate Gibbs sampler. Numerical experiments conducted on synthetic multifractal images and real data demonstrate the excellent performance of the proposed approach for image segmentation, outperforming several baselines including local multifractality-based approaches, classical texture segmentation methods as well as recent deep learning strategies. Future work includes reducing computational time through parallel implementations (e.g., GPUs) and the extension of the model to data with more than two dimensions (such as depth or spectral data volumes [belmekki20233d]), in which multidimensional/multivariate correlations can be included and exploited in the model. Other potential directions are the evaluation of the proposed approach in scenarios with corrupted spatial information, such as noisy, missing or compressed data (as in [leon2020online]) , and the integration of the proposed method with deep learning networks (as in [xiong2025fractal]).